A scientist wants to find the square of the distance d from a source to a detector. Instead of directly measuring d, she measures the distances along an x and y axis and their uncertainties. Then she uses the Pythagorean Theorem to find d2:

d2 = x2+ y2

Question 4 (1 point)

Given an uncertainty in x, what is the absolute uncertainty in x2?

Form: r=xn

Rule: Δr/r =nΔx/x

Question 4 options:

A)

Δ(x2) = 2(Δx)x

B)

Δ(x2) = 2Δx

C)

Δ(x2) = 2Δx/x
Answer #1

r=x^n

rac{Delta r}{r}=rac{nDelta x}{x}

Substitute r=x^2

rac{Delta (x^2)}{x^2}=rac{2Delta x}{x}

(2.2) 2.r.ar-2

Hence answer is old{A}),,Delta (x^2)=2(Delta x)x